Let p be the population proportion for the following condition. Find the point estimates for p and q.
A study of 4687 adults from country A found that 2754 were obese or overweight.
The point estimate for p,^p,is........
Round to 3 decimal places as needed
Can someone please help me to solve a few of these. I keep getting stuck on the same parts.Thank you very much!

Answer by jam007(5)

The point estimate of the population proportion p^ = 2754/4687 =0.588
The point estimate of q is q^ = 1-p= 1-0.588= 0.412